Find two numbers with a geometric mean of sqrt 24 given that one number is two more than the other.
Answer by xcentaur(357)

let the numbers be a and b.
then geometric mean=(a+b)/2


given second number is two greater that first,b=a+2


therefore geometric mean=(a+a+2)/2=(2a+2)/2=[2(a+1)]/2=(a+1)


From the question,it is required geometric mean be equal to sqrt24.


Then we get,
a+1=sqrt(24)
(a+1)=sqrt(2*2*2*3)
(a+1)=2sqrt(6)
a=2sqrt(6)-1


then value of b=a+2
=2sqrt(6)-1+2
=2sqrt(6)+1


cross check:
geometric mean of [2sqrt(6)-1] and [2sqrt(6)+1] is=
={[2sqrt(6)-1]+[2sqrt(6)+1]}/2
={2[2sqrt(6)]}/2
=2sqrt(6)
which is equal to sqrt(24)


Hence these numbers are correct
[2sqrt(6)-1]

[2sqrt(6)+1]


Hope this helps,
good luck.